# Limit Computation Over Posets via Minimal Initial Functors

### v1

@@ Abstract.
It is well known that limits can be computed by restricting along an initial functor, and that this can often simplify limit computation. We systematically study the algorithmic implications of this idea for diagrams indexed by a finite poset. We say an initial functor $F\colon C\to D$ with $C$ small is minimal if the sets of objects and morphisms of $C$ each have minimum cardinality, among the sources of all initial functors with target $D$. For $Q$ a finite poset or $Q\subseteq\mathbb{N}^{d}$ an interval (i.e., a convex, connected subposet), we describe all minimal initial functors $F\colon P\to Q$ and in particular, show that $F$ is always a poset inclusion. We give efficient algorithms to compute a choice of $F$. In the case that $Q\subseteq\mathbb{N}^{d}$ is an interval, we give asymptotically optimal bounds on $|P|$, the number of relations in $P$ (including identities), in terms of the number $n$ of minima of $Q$: We show that $|P|=\Theta(n)$ for $d\leq 3$, and $|P|=\Theta(n^{2})$ for $d>3$. We apply these results to give new bounds on the cost of computing $\lim G$ for a functor $G\colon Q\to\mathbf{Vec}$ valued in vector spaces. For $Q$ connected, we also give new bounds on the cost of computing the generalized rank of $G$ (i.e., the rank of the induced map $\lim G\to\operatorname{colim}G$), which is of interest in topological data analysis.
@@ 1. Introduction
Let $\mathbf{Vec}$ denote the category of vector spaces over a fixed field $\mathbbm{k}$. Given a category $D$ with finitely many objects and morphisms and a functor $G\colon D\to\mathbf{Vec}$, a standard equalizer formula [mac2013categories, Theorem V.2.2] expresses $\lim G$ as a kernel of a matrix, which can be computed via Gaussian elimination. However, this matrix can be large, and reducing it can be expensive.
@@
In this paper, we study these implications in the case that $D=Q$ is a poset and $E=\mathbf{Vec}$. (More generally, the core ideas of our approach apply when $E$ is any category where products and equalizers can be computed.) We develop theory and algorithms both for the case that $Q$ is a finite poset and the case that $Q$ is an interval in $\mathbb{N}^{d}$, (i.e., a connected, convex subposet; see Definition˜2.3). Here, $\mathbb{N}^{d}$ is given the product partial order, i.e., $(y_{1},\ldots,y_{d})\leq(z_{1},\ldots,z_{d})$ if and only if each $y_{i}\leq z_{i}$. Our motivation for considering intervals in $\mathbb{N}^{d}$ arises from computational questions in topological data analysis, and specifically in multiparameter persistent homology [carlsson2009theory, botnan2023introduction], where functors $\mathbb{N}^{d}\to\mathbf{Vec}$ are the central algebraic objects of study.
We develop a theory of minimal initial functors $F\colon P\to Q$, i.e., those for which the sets of objects and morphisms of $P$ both have minimum cardinality, among all sources of all initial functors. We give algorithms for computing such functors, and work out the consequences of these results for limit computation. Here, we summarize our main results, deferring the precise statements to Section˜3.
@@
    We present a structure theorem for minimal initial functors, Theorem˜3.5, which establishes the existence of a minimal initial functor $F\colon P\to Q$ and gives an explicit description of all such $F$. Specifically, the theorem says that, up to unique isomorphism, any such $F$ is the inclusion of a certain type of subposet $P\subseteq Q$, which we call an initial scaffold. While $Q$ can have multiple non-isomorphic initial scaffolds, all have the same underlying set of objects (but different sets of relations). As we discuss in Remark˜3.7, related results concerning full subposets have appeared in a recent paper of Brüstle, Desrochers, and Leblanc [brustle2025generalized].
@@
    For $Q\subseteq\mathbb{N}^{d}$ an interval, Theorem˜3.10 bounds $|P|$, the number of relations of $P$ (including identity relations), in terms of the number $n$ of minima of $Q$, as follows:
@@
    To obtain these bounds, we establish a connection between initial scaffolds of intervals in $\mathbb{N}^{d}$ and the support of Betti numbers of monomial ideals. We then apply a well-known bound on the Betti numbers of monomial ideals, due to Bayer, Peeva, and Sturmfels [bayer1998monomial].
@@
    For $Q$ any finite poset, we give an algorithm to compute a choice of $F$ from the Hasse diagram of $Q$. We also give two specialized algorithms to compute $F$ when $Q\subseteq\mathbb{N}^{d}$ is an interval, one for the case $d\leq 3$ and one for arbitrary $d$. Theorems˜3.13 and 3.17 bound the complexity of each algorithm. Our algorithms for the interval case do not require the Hasse diagram of $Q$ as input, but instead take a more parsimonious representation of $Q$, which we call the upset presentation of $Q$; see Definition˜3.14. If the interval $Q\subseteq\mathbb{N}^{d}$ is finite, then we can instead specify $Q$ via its sets of minima and maxima.
@@
    We apply these results to derive new algorithms and bounds for computing the (co)limit of a functor $G\colon Q\to\mathbf{Vec}$ (Corollaries˜3.23 and 3.24). The statements of these bounds depend on how $G$ is input to the algorithms. For one concrete corollary of these bounds, suppose $Q\subseteq\mathbb{N}^{d}$ is a finite interval with a total of $s$ minima and maxima and we are given a free presentation $F_{1}\to F_{0}$ of $G$ where $F_{1}$ and $F_{0}$ have total rank $r$. Letting $\omega<2.373$ denote the exponent of matrix multiplication, Corollary˜3.31 implies that the cost of computing $\lim G$ is
    $O(s\log s+r^{3})$  $\quad\textup{ for }d=2,$
    $O((rs)^{\omega})$  $\quad\textup{ for }d=3,$
    $O(r^{\omega}s^{\omega+1})$  $\quad\textup{ for }d>3.$
@@
    Given a connected, finite poset $Q$ and functor $G\colon Q\to\mathbf{Vec}$, we apply our results to bound the cost of computing the generalized rank of $G$, which is the rank of the induced map $\lim G\to\operatorname{colim}G$; see Corollaries˜3.36 and 3.37. As we discuss in Section˜3.5, this problem is of interest in TDA and has been studied in prior work [DKM22, dey2024computing, asashiba2024interval]. Our bounds extend and improve on prior bounds by Dey, Kim, and Mémoli [DKM22] and by Dey and Xin [dey2024computing].
@@
In Section˜2, we cover preliminaries about posets, limits, initial functors, and $\mathbf{Vec}$-valued functors. Section˜3 is the heart of our narrative: Here, we precisely state the main definitions and results of our work, together with examples and context. The remaining sections of the paper give the algorithms and proofs underlying our main results: Section˜4 gives the proof of Theorem˜3.5, our structure theorem for minimal initial functors; Section˜5 gives the proof of our size bound Theorem˜3.10; Section˜6 presents our algorithms for computing initial scaffolds, and proves the bounds on their complexity; Section˜7 completes the proofs of our bounds on limit computation; and Section˜8 proves our bounds on generalized rank computation. Section˜9 briefly discussions directions for future work.
@@ Acknowledgments
We thank Ezra Miller for pointing out that Lemma˜5.12 follows from a version of Hochster’s formula [miller2005combinatorial, Theorem 1.34]. TD acknowledges the support of NSF grants DMS 2301360, and CCF 2437030. ML acknowledges the support of a grant from the Simons Foundation (Award ID 963845).
@@ 2.1. Categories and Posets
We assume familiarity with a few very basic notions in category theory, e.g., categories, (full) functors, and natural transformations. For an introduction to category theory, see, e.g., the textbook [riehl2017category, mac2013categories].
We denote the collections of objects and morphisms of a category $C$ as $\operatorname{Ob}C$ and $\hom C$, respectively. For $c,c^{\prime}\in\operatorname{Ob}C$, the set of morphisms from $c$ to $c^{\prime}$ is denoted $\hom(c,c^{\prime})$. We use subscripts to denote the action of a functor on objects and morphisms. That is, given a functor $F\colon C\to D$ and $c\in\operatorname{Ob}C$ we let $F_{c}=F(c)$, and given a morphism $g\colon c\to c^{\prime}$, we let $F_{g}=F(g)$. We call $C$ thin if $|\hom(c,c^{\prime})|\leq 1$ for all $c,c^{\prime}\in\operatorname{Ob}C$. If $C$ is thin, then $g\colon c\to c^{\prime}$ is determined by its source and target, so we write $F_{cc^{\prime}}=F(g)$.
@@
When $p\leq q$ and $p\neq q$, we write $p<q$. We often abuse notation slightly and let $Q$ denote the poset $(Q,\leq)$. By a further slight abuse, the elements of $\leq$ are called relations.
We regard a poset $Q$ as a thin category where $\operatorname{Ob}Q$ and $\hom Q$ are the sets of elements and relations of $Q$, respectively. A poset $P$ is called a subposet of $Q$ if $\operatorname{Ob}P\subseteq\operatorname{Ob}Q$ and $\hom P\subseteq\hom Q$. In this case, we write $P\subseteq Q$. We call $P$ a full subposet if the inclusion functor $P\hookrightarrow Q$ is full, i.e., $p\leq p^{\prime}$ in $P$ whenever $p,p^{\prime}\in\operatorname{Ob}P$ and $p\leq p^{\prime}$ in $Q$.
Intersections of posets are defined in the obvious way, i.e., by taking intersections of both the underlying sets and the sets of relations. We define the union of posets analogously. In general a union of posets needn’t be a poset, as it may fail the transitivity property, but all unions considered in this paper satisfy transitivity.
@@
We call a category $C$ connected if $C$ is non-empty (i.e., has at least one object), and for each pair of objects $c,c^{\prime}\in\operatorname{Ob}C$, there exists a finite zigzag of morphisms
[…]
connecting $c$ and $c^{\prime}$. We call such a zigzag a path from $c$ to $c^{\prime}$. A component of $C$ is a maximal full, connected subcategory of $C$.
Note that a finite poset $Q$ is connected if and only if the undirected graph underlying its Hasse diagram is a connected graph.
@@
    For posets $P\subseteq Q$, the (closed) upset of $P$ in $Q$ is the set
    $\operatorname{up}[P,Q]\coloneqq\{q\in Q\mid q\geq p\textup{ for some }p\in P\}.$
@@
When $Q$ is clear from context, we sometimes omit it from the above notation, i.e.,
$\operatorname{dn}(q)$ $\coloneqq\operatorname{dn}(q,Q),$
@@ 2.2. Limits
Given a small category $C$, a category $D$, and a functor $G\colon C\to D$, a cone on $G$ consists of an object $d\in\operatorname{Ob}D$, together with a morphism $\gamma_{c}\colon d\to G_{c}$ for each $c\in\operatorname{Ob}C$, called a cone map, such that for all morphisms $g\colon c\to c^{\prime}$ in $\hom C$, we have $\gamma_{c^{\prime}}=G_{g}\circ\gamma_{c}$. We write the cone as $(d,\gamma)$.
@@
The limit of $G$ is a cone $(d,\gamma)$ such that for any other cone $(d^{\prime},\gamma^{\prime})$, we have a unique morphism $\kappa\colon d^{\prime}\to d$ with $\gamma^{\prime}_{c}=\gamma_{c}\circ\kappa$ for all $c\in\operatorname{Ob}C$.
When they exist, limits are unique up to unique isomorphism, which justifies speaking of the limit of $G$. If $(d,\gamma)=\lim G$, we sometimes abuse terminology slightly and refer to $d$ as $\lim G$, particularly when $\gamma$ is clear from context. Cocones and colimits are defined dually.
Given functors $F\colon C\to D$ and $G\colon D\to E$ such that $\lim G$ and $\lim(G\circ F)$ both exist, the limit cone $(\lim G,\gamma)$ on $G$ restricts to a cone $(\lim G,\kappa)$ on $G\circ F$. The universality of $\lim(G\circ F)$ then yields a morphism $\lim G\to\lim(G\circ F)$ which commutes with the relevant cone maps. This is functorial in the sense that, given another functor $F^{\prime}\colon C^{\prime}\to C$, the morphism $\lim G\to\lim(G\circ F\circ F^{\prime})$ factors as
$\lim G\to\lim(G\circ F)\to\lim(G\circ F\circ F^{\prime}).$

### v2

@@ Abstract.
It is well known that limits can be computed by restricting along an initial functor, and that this often simplifies limit computation. We systematically study the algorithmic implications of this idea for diagrams indexed by a finite poset. We say an initial functor $F\colon C\to D$ with $C$ small is minimal if the sets of objects and morphisms of $C$ each have minimum cardinality, among the sources of all initial functors with target $D$. For $Q$ a finite poset or $Q\subseteq\mathbb{N}^{d}$ an interval (i.e., a convex, connected subposet), we describe all minimal initial functors $F\colon P\to Q$ and in particular, show that $F$ is always a subposet inclusion. We give efficient algorithms to compute a choice of minimal initial functor. In the case that $Q\subseteq\mathbb{N}^{d}$ is an interval, we give asymptotically optimal bounds on $|P|$, the number of relations in $P$ (including identities), in terms of the number $n$ of minima of $Q$: We show that $|P|=\Theta(n)$ for $d\leq 3$, and $|P|=\Theta(n^{2})$ for $d>3$. We apply these results to give new bounds on the cost of computing $\lim G$ for a functor $G\colon Q\to\mathbf{Vec}$ valued in vector spaces. For $Q$ connected, we also give new bounds on the cost of computing the generalized rank of $G$ (i.e., the rank of the induced map $\lim G\to\operatorname{colim}G$), which is of interest in topological data analysis.
@@ 1. Introduction
Let $\mathbf{Vec}$ denote the category of vector spaces over a fixed field $\mathbbm{k}$. Given a category $D$ with finitely many morphisms and a pointwise finite-dimensional functor $G\colon D\to\mathbf{Vec}$, a standard equalizer formula [33, Theorem V.2.2] expresses $\lim G$ as the kernel of a matrix, which can be computed via Gaussian elimination. However, if $D$ has many objects and morphisms and the vector spaces of $G$ have large dimensions, then this matrix will be large and computing its kernel can be costly.
@@
In this paper, we study these implications in the case that $D=Q$ is a poset and $E=\mathbf{Vec}$. (More generally, the core ideas of our approach apply when $E$ is any category where products and equalizers can be computed.) We develop theory and algorithms both for the case that $Q$ is a finite poset and the case that $Q$ is an interval in $\mathbb{N}^{d}$ (i.e., a connected, convex subposet; see Definition˜2.4). Here, $\mathbb{N}^{d}$ is given the product partial order, i.e., $(y_{1},\ldots,y_{d})\leq(z_{1},\ldots,z_{d})$ if and only if each $y_{i}\leq z_{i}$. Our motivation for considering intervals in $\mathbb{N}^{d}$ arises from computational questions in topological data analysis, and specifically in multiparameter persistent homology [10, 7], where functors $\mathbb{N}^{d}\to\mathbf{Vec}$ are the central algebraic objects of study.
We develop a theory of minimal initial functors $F\colon P\to Q$, i.e., those for which the sets of objects and morphisms of $P$ both have minimum cardinality, among the sources of all initial functors with target $Q$. We give algorithms for computing such functors and apply them to the problem of limit computation. Here, we summarize our main results, deferring the precise statements to Section˜3.
@@
    We present a structure theorem for minimal initial functors, Theorem˜3.5, which establishes the existence of a minimal initial functor $F\colon P\to Q$ and gives an explicit description of all such $F$. Specifically, the theorem says that, up to canonical isomorphism, any such $F$ is the inclusion of a certain type of subposet $P\subseteq Q$, which we call an initial scaffold. While $Q$ can have multiple non-isomorphic initial scaffolds, all have the same underlying set of objects (but different sets of relations). As we discuss in Remark˜3.7, related results concerning full subposets have appeared in a recent paper of Brüstle, Desrochers, and Leblanc [8].
@@
    For $Q\subseteq\mathbb{N}^{d}$ an interval, Theorem˜3.10 bounds $|P|$, the number of relations of an initial scaffold $P\subseteq Q$ (including identities), in terms of the number $n$ of minima of $Q$, as follows:
@@
    To obtain these bounds, we establish a connection between initial scaffolds of intervals in $\mathbb{N}^{d}$ and the support of Betti numbers of monomial ideals. We then apply a well-known bound on the Betti numbers of monomial ideals, due to Bayer, Peeva, and Sturmfels [5].
@@
    For $Q$ any finite poset, we give an algorithm to compute a choice of initial scaffold $P\subseteq Q$ from the Hasse diagram of $Q$. We also give two specialized algorithms to compute $P$ when $Q\subseteq\mathbb{N}^{d}$ is an interval, one for the case $d\leq 3$ and one for arbitrary $d$. Theorems˜3.13 and 3.16 bound the complexity of each algorithm. Our algorithms for the interval case do not require the Hasse diagram of $Q$ as input, but instead take a more parsimonious representation of $Q$, which we call the upset presentation of $Q$; see Definition˜3.14. If the interval $Q\subseteq\mathbb{N}^{d}$ is finite, then we can instead specify $Q$ via its sets of minima and maxima.
@@
    We apply these results to derive new algorithms and bounds for computing the (co)limit of a functor $G\colon Q\to\mathbf{Vec}$ (Corollaries˜3.22 and 3.23). The statements of these bounds depend on how $G$ is input to the algorithms. For one concrete corollary of these bounds, suppose $Q\subseteq\mathbb{N}^{d}$ is a finite interval with a total of $n$ minima and maxima and we are given a free presentation $F_{1}\to F_{0}$ of $G$ where $F_{1}$ and $F_{0}$ have total rank $r$. Letting $\omega<2.373$ denote the exponent of matrix multiplication, Corollary˜3.30 implies that the cost of computing $\lim G$ is
    $O(n\log n+r^{3})$  $\quad\textup{ for }d=2,$
    $O((nr)^{\omega})$  $\quad\textup{ for }d=3,$
    $O(n^{4}+n^{\omega+1}r^{\omega})$  $\quad\textup{ for }d>3.$
@@
    Given a connected, finite poset $Q$ and functor $G\colon Q\to\mathbf{Vec}$, we apply our results to bound the cost of computing the generalized rank of $G$, i.e., the rank of the induced map $\lim G\to\operatorname{colim}G$; see Corollaries˜3.35 and 3.36. As we discuss in Section˜3.5, this problem is of interest in TDA and has been studied in prior work [16, 17, 2]. Our bounds extend and improve on prior bounds by Dey, Kim, and Mémoli [16] and by Dey and Xin [17].
@@
Section˜2 covers preliminaries. Section˜3 is the backbone of the paper: Here, we precisely state the main definitions and results of our work, together with examples and context. The remaining sections of the paper give the algorithms and proofs underlying our main results: Section˜4 gives the proof of Theorem˜3.5, our structure theorem for minimal initial functors; Section˜5 gives the proof of our size bound Theorem˜3.10; Section˜6 presents our algorithms for computing initial scaffolds, and proves the bounds on their complexity; Section˜7 completes the proofs of our bounds on limit computation; and Section˜8 proves our bounds on generalized rank computation. Section˜9 briefly discusses directions for future work.
@@ Acknowledgments
We thank Ezra Miller for pointing out that Lemma˜5.12 follows from a version of Hochster’s formula [35, Theorem 1.34]. TD acknowledges the support of NSF grants DMS 2301360, and CCF 2437030. ML acknowledges the support of a grant from the Simons Foundation (Award ID 963845).
@@ 2. Preliminaries
In this section, we review background on categories, posets, limits, initial functors, and $\mathbf{Vec}$-valued functors that we will need throughout the paper. While we have aimed for an accessible treatment of the necessary category theory, we do assume familiarity with a few very basic notions, e.g., functors and natural transformations. For background on category theory, see, e.g., the textbooks [39, 33].
@@ 2.1. Categories and Posets
We denote the collections of objects and morphisms of a category $C$ as $\operatorname{Ob}C$ and $\hom C$, respectively. For $c,c^{\prime}\in\operatorname{Ob}C$, the set of morphisms from $c$ to $c^{\prime}$ is denoted $\hom(c,c^{\prime})$. We use subscripts to denote the action of a functor on objects and morphisms. That is, given a functor $F\colon C\to D$ and $c\in\operatorname{Ob}C$ we let $F_{c}=F(c)$, and given a morphism $\gamma\colon c\to c^{\prime}$, we let $F_{\gamma}=F(\gamma)$. We call $C$ thin if $|\hom(c,c^{\prime})|\leq 1$ for all $c,c^{\prime}\in\operatorname{Ob}C$. If $C$ is thin, then $\gamma\colon c\to c^{\prime}$ is determined by its source and target, so we write $F_{cc^{\prime}}=F(\gamma)$.
Recall that a functor $F\colon C\to D$ is called faithful (respectively, full) if for each $c,c^{\prime}\in\operatorname{Ob}C$, the induced map $\hom(c,c^{\prime})\to\hom(F_{c},F_{c^{\prime}})$ is an injection (respectively, surjection).
Definition 2.1.
An embedding is a faithful functor that is injective on objects.
@@
When $p\leq q$ and $p\neq q$, we write $p<q$. We often abuse notation slightly and let $Q$ denote the poset $(Q,\leq)$. By a further slight abuse, we refer to the elements of $\leq$ as relations.
We regard a poset $Q$ as a thin category, where $\operatorname{Ob}Q$ and $\hom Q$ are the sets of elements and relations of $Q$, respectively. A poset $P$ is called a subposet of $Q$ if $\operatorname{Ob}P\subseteq\operatorname{Ob}Q$ and $\hom P\subseteq\hom Q$. In this case, we write $P\subseteq Q$. By a slight abuse of terminology, we sometimes also refer to the inclusion map $P\hookrightarrow Q$ as a subposet. We call $P$ a full subposet if the inclusion functor $P\hookrightarrow Q$ is full, i.e., $p\leq p^{\prime}$ in $P$ whenever $p,p^{\prime}\in\operatorname{Ob}P$ and $p\leq p^{\prime}$ in $Q$.
Intersections of posets are defined in the obvious way, i.e., by taking intersections of both the underlying sets and the sets of relations. We define the union of posets analogously. In general, a union of posets need not be a poset, as it may fail the transitivity property, but all unions considered in this paper satisfy transitivity and thus are posets.
@@
Let $C$ be a category.
  • (i)
    For $c,c^{\prime}\in\operatorname{Ob}C$, a path from $c$ to $c^{\prime}$ in $C$ is a sequence
    $c=c_{1},\ldots,c_{k}=c^{\prime}$
    in $\operatorname{Ob}C$ such that either $\hom(c_{i},c_{i+1})$ or $\hom(c_{i+1},c_{i})$ is nonempty for each $i\in\{1,\ldots,k-1\}$.
  • (ii)
    $C$ is connected if $C$ is non-empty (i.e., has at least one object) and there exists a path between each pair of objects of $C$.
  • (iii)
    A component of $C$ is a maximal connected subcategory of $C$.
Note that a finite poset $Q$ is connected if and only if the undirected graph underlying its Hasse diagram is a connected graph. Note also that each object or morphism of $C$ belongs to exactly one component of $C$.
@@
Let $\mathsf{Z}$ be the full subposet of $\mathbb{Z}^{2}$ with the following elements:
$\{(z,z)\mid z\in\mathbb{Z}\}\cup\{(z,z+1)\mid z\in\mathbb{Z}\}.$
A zigzag poset is a poset isomorphic to a finite interval in $\mathsf{Z}$.
Example 2.6.
The poset $P$ with the following Hasse diagram is a zigzag poset.
v{v}w{w}x{x}y{y}z,{z,}
Definition 2.7.
@@
    generalizing (i), the downset of $Q^{\prime}\subseteq Q$ is the set
    $\operatorname{dn}[Q^{\prime},Q]\coloneqq\{p\in Q\mid p\leq q\textup{ for some }q\in Q^{\prime}\},$
  • (iv)
    dually, the upset of $Q^{\prime}\subseteq Q$ is the set
    $\operatorname{up}[Q^{\prime},Q]\coloneqq\{p\in Q\mid p\geq q\textup{ for some }q\in Q^{\prime}\}.$
@@
When $Q$ is clear from context, we sometimes omit it from the above notation, e.g.,
@@ 2.2. Limits
Given a small category $C$, a category $D$, and a functor $G\colon C\to D$, a cone on $G$ consists of an object $d\in\operatorname{Ob}D$, together with a morphism $\delta_{c}\colon d\to G_{c}$ for each $c\in\operatorname{Ob}C$, called a cone map, such that for all morphisms $\gamma\colon c\to c^{\prime}$ in $\hom C$, we have $\delta_{c^{\prime}}=G_{\gamma}\circ\delta_{c}$. We write the cone as $(d,\delta)$.
Example 2.8.
If $P$ is the poset of Example˜2.6, then a cone on a functor $G\colon P\to D$ is a commutative diagram in $D$ of the following shape, extending $G$:
[…]
@@
The limit of $G$ is a cone $(d,\delta)$ such that for any other cone $(d^{\prime},\delta^{\prime})$, we have a unique morphism $\kappa\colon d^{\prime}\to d$ satisfying $\delta^{\prime}_{c}=\delta_{c}\circ\kappa$ for all $c\in\operatorname{Ob}C$.
When they exist, limits are unique up to unique isomorphism, which justifies speaking of the limit of $G$. If $(d,\delta)=\lim G$, we sometimes abuse terminology slightly and refer to $d$ as $\lim G$, particularly when $\delta$ is clear from context. Cocones and colimits are defined dually.
@@
Given functors $F\colon C\to D$ and $G\colon D\to E$ such that $\lim G$ and $\lim(G\circ F)$ both exist, the limit cone $(\lim G,\delta)$ on $G$ restricts to a cone $(\lim G,\delta^{\prime})$ on $G\circ F$. The universality of $\lim(G\circ F)$ then yields a morphism $\lim G\to\lim(G\circ F)$ such that the following triangle commutes for each $c\in\operatorname{Ob}C$, where the diagonal arrows are the limit cone maps:
$\lim G$$\lim(G\circ F)$$G_{F_{c}}=(G\circ F)_{c}$}
This is functorial in the sense that, given another functor $F^{\prime}\colon C^{\prime}\to C$, the maps on limits form the following commutative triangle:
$\lim G$$\lim(G\circ F)$$\lim(G\circ F\circ F^{\prime})$}